surface area of cubiod with top side open and its length =150cm,breadth=125cm and height =65cm
Answer by Cromlix(4381)

Surface Area :=
2 * 150cm * 65cm = 19500cm^2
2 * 125cm * 65cm = 16250cm^2
1 * 150cm * 125cm = 18750cm^2
Total = 54500cm^2
Hope this helps.
:-)